# Coordinate In and Value Out: Training Flow Transformers in Ambient Space

### v1

@@ 5 Conclusion
We introduced Ambient Space Flow Transformers (ASFT), a flow matching generative model designed to operate directly in ambient space. Our approach dispenses with the practical complexities of training latent space generative models, such as the dependence on domain-specific compressors for different data domains or tuning of hyper-parameters of the data compressor (i.e. adversarial weight, KL term, etc.). We introduced a conditionally independent point-wise training objective that decomposes the target vector field and allows to continuously evaluate the generated samples, enabling resolution changes at inference time. This training objective also improves training efficiency since it allows us to sub-sample the target vector field during training. Our results on both image and 3D point cloud benchmarks show the strong performance of ASFT as well as its trivial adaption across modalities. In conclusion, ASFT represents a promising direction for flow matching generative models, offering a powerful and domain-agnostic framework. Future work could explore further improvements in training efficiency and investigate co-training of multiple data domains to enable multi-modality generation in an end-to-end learning paradigm.
Acknowledgements
We thank Jiatao Gu for helpful discussions and Jen-Hao Rick Chang for helping set up experiments on Objaverse.
@@ Appendix A Model Configuration and Training Settings
We provide detailed model configurations and training settings of ASFT for image (Tab. 6) and point cloud (Tab. 7) generation. For image generation, we develop model sizes small (S), base (B), large (L), and extra large (XL) to approximately match the number of parameters in previous works (Peebles & Xie, 2023). Similarly, for point cloud generation, we train a base sized model roughly matching the number of parameters in LION (Vahdat et al., 2022) (i.e. 110M parameters), and a ASFT-L which contains about twice the number of parameters as ASFT-B. For image experiments we implement the “psuedo” coordinate of latents as 2D grids and coordinate-value pairs are assigned to different latents based on their distances to the latent coordinates. Whereas in point cloud generation, since calculating the pair-wise distances in 3D space can be time consuming, we assign input elements to latents through a hash code, so that neighboring input elements are likely (but not certainly) to be assigned to the same latent token. We found that the improvements of spatial aware latents in 3D to not be as substantial as in the 2D image setting, so we report results with a vanilla PerceiverIO architecture for simplicity. To embed coordinates, we apply standard Fourier positional embedding (Vaswani et al., 2017) for ambient space coordinate input in both encoder and decoder. The Fourier positional embedding is also applied to the “psuedo” coordinate of latents. On image generation, we found that applying rotary positional embedding (RoPE) (Su et al., 2024) slightly improves the performance of ASFT. Therefore, RoPE is employed for largest ASFT-XL model. For all the models including image and 3D point cloud experiments, we share the following training parameters except the training_steps across different experiments. On image generation, all models are trained with batch size 256, except for ASFT-XL reported in Tab. 2 and Tab. 3, which are trained for 1.7M steps with batch size 512. On ShapeNet, ASFT models are trained for 800K iterations with a batch size of 16. On Objaverse, we follow the configuration ASFT-XL listed in Tab. 7 and the model is trained with batch size 384 for 500k iterations.
@@
In Tab. 8, we also compare the size of models trained on ImageNet-256, training cost (i.e. product of batch size and training iterations), and inference cost (i.e. NFE, number of function evaluation). Note that for models that achieve better performance than ASFT, many of them are trained for more iterations. In addition, at inference time ASFT applies simple first order Euler sampler with 100 sampling steps, which uses less NFE than many other baselines.
@@
ASFT-S 12 384 1024 6 1 35M
ASFT-B 12 768 1024 12 1 138M
ASFT-L 24 1024 1024 16 1 458M
ASFT-XL 28 1152 1024 16 2 733M
Table 6: Detailed configurations of ASFT for image generation.
Model Layers Hidden size #Latents Heads Decoder layers #Params
ASFT-B 9 512 1024 4 1 108M
ASFT-L 12 512 1024 4 1 204M
ASFT-XL 28 1152 1024 16 1 866M
Table 7: Detailed configurations of ASFT for point cloud generation.
@@
Table 8: Comparison of ASFT and baselines in # params and training cost (i.e. product of batch size and training iterations). Some numbers are borrowed from Crowson et al. (2024).
Appendix B Performance vs Training Compute
@@
Figure 5: Comparing the performance vs total training compute comparison of ASFT and DiT (Peebles & Xie, 2023).
@@
We compare the performance vs total training compute of ASFT and DiT (Peebles & Xie, 2023) in Gflops. ASFT-linear denotes the variant of ASFT where the cross-attention in the spatial aware encoder is replaced with grouping followed by a linear layer. We found this could be an efficient variant of standard ASFT while still achieving competitive performance. Fig. 5 shows the comparison of the training compute in Gflops vs FID-50K between ASFT and latent diffusion model DiT (Peebles & Xie, 2023) including the tranining compute of the first stage VAE. We estimate the training cost of VAE based the model card listed in HuggingFace [^2] https://huggingface.co/stabilityai/sd-vae-ft-mse. As shown, the training cost of VAE is not negligible and reasonable models with FID $\approx 6.5$ can be trained for the same cost.
@@
Admittedly, under equivalent training Gflops, ASFT achieves comparable but not as good performance as DiT in terms of FID score (with a difference smaller than $1.65$ FID points). We attribute this gap to the fact that DiT’s VAE was trained on a dataset much larger than ImageNet, using a domain-specific architecture (e.g. a convolutional U-Net). We believe that the simplicity of implementing and training ASFT models in practice, and the trivial extension to different data domains (as shown in Sect. 4.3) are strong arguments to counter an FID difference of smaller than $1.65$ points.
Appendix C Architecture Ablation
We also provide an architecture ablation in Tab. 9 showcasing different design decisions. We compare two variants of Transformer-based architectures ASFT: a vanilla PerceiverIO that directly operates on ambient space, but without using spatial aware latents and ASFT. As it can be seen, the spatially aware latents introduced in ASFT greatly improve performance across all metrics in the image domain, justifying our design decisions. We note that we did not observe the same large benefits for 3D point clouds, which we hypothesize can be due to their irregular structure.
@@ Appendix C Architecture Ablation
Table 9: Benchmarking vanilla PerceiverIO and ASFT with spatially aware latents on LSUN-Church-256 (Yu et al., 2015).
Appendix D Resolution Agnostic Generation: Quantitative Analysis
Due to the fact that ASFT decodes each coordinate-value pair independently given ${\bm{z}}_{f_{t}}$, during inference time one can decode as many coordinate-value pairs as desired, therefore allowing resolution to change during inference. We now quantitatively evaluate the performance of ASFT in this setting. In Tab. 10 we compare the FID of different recipes. First, ASFT is trained on FFHQ-256 and bilinear or bicubic interpolation is applied to generated samples to get images at 512. On the other hand, ASFT can directly generate images at resolution 512 by simply increasing the number of coordinate-value pairs during inference without further tuning. As shown in Tab. 10 , ASFT achieves lower FID when compared with other manually designed interpolation methods, showcasing the benefit of developing generative models on ambient space.
ASFT Bilinear Bicubic
FID($\downarrow$) 23.09 35.05 24.34
Table 10: FID of different super resolution methods to generate images at resolution $512\times 512$ for ASFT trained on FFHQ-256.
Appendix E Additional ImageNet Samples
@@
We show uncurated samples of different classes from ASFT-XL trained on ImageNet-256 in Fig. 6 and Fig. 7. Guidance scales in CFG are set as $4.0$ for loggerhead turtle, macaw, otter, coral reef and $2.0$ otherwise.
@@
Figure 6: Uncurated samples of class labels: loggerhead turtle (33), macaw (88), golden retriever (207), otter (360) and red panda (387), and panda (388) from ASFT trained on ImageNet-256.
@@
Figure 7: Uncurated samples of class labels: palace (698), space shuttle (812), ice cream (928), pizza (963), coral reef (973), and valley (979) from ASFT trained on ImageNet-256.
Appendix F Additional ShapeNet Samples
We show uncurated samples from ASFT-L trained jointly on 55 ShapeNet categories in Fig. 8.
@@ Appendix F Additional ShapeNet Samples
Appendix G Additional Objaverse Samples
We show conditional images and generated samples from ASFT trained on Objaverse in Fig. 9 and 10.
Figure 9: Image-conditioned point clouds with $16k$, $32k$, $64k$, and $128k$ points generated from an ASFT trained on Objaverse (training with $16k$ points, CFG scale 5.0).
Figure 10: Image-conditioned point clouds with $16k$, $32k$, $64k$, and $128k$ points generated from an ASFT trained on Objaverse (training with $16k$ points, CFG scale 5.0).

### v2

@@ 5 Conclusion
We introduced INRFlow, a flow matching generative model for continuous function spaces designed to operate directly in ambient space. Our approach dispenses with the practical complexities of training latent space generative models (Du et al., 2021a; Dupont et al., 2022a), such as the dependence on domain-specific compressors for different data domains or tuning of hyper-parameters of the data compressor (i.e. adversarial weight, KL term, etc.). Inspired by deterministic encoding of INRs, we introduced a conditionally independent point-wise training objective that decomposes the target vector field and allows to continuously evaluate the generated samples, similar to INRs, enabling resolution changes at inference time.
Our results on both image, 3D point cloud and protein folding benchmarks show the strong performance of INRFlow, as well as, its trivial adaption across modalities which we believe is the cornerstone of a good generative modeling architecture. In conclusion, INRFlow represents a promising direction for flow matching generative models, offering a powerful and domain-agnostic framework. Future work could explore further improvements in training efficiency applying tricks orthogonal to our contribution (Sehwag et al., 2024) and investigate co-training of multiple data domains to enable multi-modality generation in an end-to-end learning paradigm.
Impact Statement
This paper concerns the generative modeling methodology. While we do not see immediate societal implications from our technical contribution, there are potential consequences when it is used in training foundational generative models.
@@ Appendix A Model Configuration and Training Settings
A.1 Image
We provide detailed model configurations and training settings of INRFlow for image generation in Tab. 6. We develop model sizes small (S), base (B), large (L), and extra large (XL) to approximately match the number of parameters in DiT (Peebles & Xie, 2023). For image experiments we implement the “psuedo” coordinate of latents as 2D grids and coordinate-value pairs are assigned to different latents based on their distances to the latent coordinates. To embed coordinates, we apply standard Fourier positional embedding (Vaswani et al., 2017) for ambient space coordinate input in both encoder and decoder. The Fourier positional embedding is also applied to the “psuedo” coordinate of latents. On image generation, we found that applying rotary positional embedding (RoPE) (Su et al., 2024) slightly improves the performance of INRFlow. Therefore, RoPE is employed for largest INRFlow-XL model. For all the models we share the following training parameters except the training_steps across different experiments. On image generation, all models are trained with batch size 256, except for INRFlow-XL reported in Tab. 2 and Tab. 3, which are trained for 1.7M steps with batch size 512.
@@ A.1 Image
In Tab. 7, we also compare the size of models trained on ImageNet-256, training cost (i.e. product of batch size and training iterations), and inference cost (i.e. NFE, number of function evaluation). Note that for models that achieve better performance than INRFlow, many of them are trained for more iterations. In addition, at inference time INRFlow applies simple first order Euler sampler with 100 sampling steps, which uses less NFE than many other baselines.
@@
INRFlow-S 12 384 1024 6 1 35M
INRFlow-B 12 768 1024 12 1 138M
INRFlow-L 24 1024 1024 16 1 458M
INRFlow-XL 28 1152 1024 16 2 733M
Table 6: Detailed configurations of INRFlow for image generation.
@@
Table 7: Comparison of INRFlow and baselines in # params and training cost (i.e. product of batch size and training iterations). Some numbers are borrowed from (Crowson et al., 2024).
A.2 Image-to-3D
For image-to-3D point cloud generation we trained INRFlow on Objaverse (Deitke et al., 2023), which contains 800k 3D objects of great variety. In particular, conditional information (i.e., an image) is integrated to our model through cross-attention. For each object in Objaverse, we sample point cloud with $16k$ points. To get images for conditioning, each object is rendered with 40 degrees field of view, $448\times 448$ resolution, at 3.5 units on the opposite sides of $x$ and $z$ axes looking at the origin. We extract features via DINOv2 (Oquab et al., 2023) which is concatenated with Plucker ray embedding (Plucker, 2018) of each patch in DINOv2 feature. In each block, the learnable latent vector $z_{f_{t}}$ cross attends to image feature. During training, the image conditioning is dropped randomly with 10% probability. Therefore, our model can also benefit from popular classifier-free guidance (CFG) to increase the guidance strength. The model is trained with batch size 384 for 500k iterations. During sampling, we use an Euler-Maruyama sampler with 500 steps to generate point clouds. We train an INRFlow-XL size model of 866M parameters similar to the one reported in Tab. 11.
One particularity for image-to-3D point cloud generation is that we assign input elements to latents through a hash code, so that neighboring input elements are likely (but not certainly) to be assigned to the same latent token. We found that the improvements of spatial aware latents in 3D to not be as substantial as in the 2D image setting, so we report results with a vanilla PerceiverIO architecture for simplicity. To embed coordinates, we apply standard Fourier positional embedding (Vaswani et al., 2017) for ambient space coordinate input in both encoder and decoder.
A.3 Protein Folding
In our experiments we use SwissProt set (Boeckmann et al., 2003) taking the ground truth structures from the AlphaFold Database (Varadi et al., 2022). We select a random set of 10k protein structures to train INRFlow. In this setting, the coordinate-value pairs represent atoms in the protein, where the "coordinate" part is a set of features of that particular atom. In particular, we use the atom features shown in Tab. 8 together with the eigenvectors of the graph laplacian for each residue, as previously done in (Wang et al., ). In particular, we found it beneficial to also concatenate to the atomic features the amino-acid embedding of its corresponding amino-acid, which we obtain from a pre-trained ESM model (ESM-650M), a language model for protein sequences masking only on sequence data and not 3D structures.
We aggregate information into spatially-aware latents by cross-attending all atoms belonging to a particular amino-acid with its particular latent. In practice, since the number of atoms for each aminoacid type is always fixed, one might also simply use a linear layer. Finally, for this task we train a XL size model for 100k iterations with batch size 256.
Name Description Range
atomic Atom type one-hot of 35 elements in dataset
degree Number of bonded neighbors $\{x:0\leq x\leq 6,x\in\mathbbm{Z}\}$
charge Formal charge of atom $\{x:-1\leq x\leq 1,x\in\mathbbm{Z}\}$
valence Implicit valence of atom $\{x:0\leq x\leq 6,x\in\mathbbm{Z}\}$
hybrization Hybrization type {sp, sp2, sp3, sp3d, sp3d2, other}
aromatic Whether on a aromatic ring {True, False}
num_rings number of rings atom is in $\{x:0\leq x\leq 3,x\in\mathbbm{Z}\}$
ESM embedding Amino-acid embedding $x\in\mathbb{R}^{1280}$
Table 8: Atomic features included in INRFlow for protein folding.
Appendix B Performance vs Model Size
To demonstrate the scalability of INRFlow we train models of different sizes including small (S), base (B), large (L), and extra-large (XL) on ImageNet-256. We show the performance of different model sizes using FID-50K in Fig. 5(a). We observe a clear improving trend when increasing the number of parameters as well as increasing training steps. This demonstrates that scaling the total training Gflops is important to improved generative results as in other ViT-based generative models (Peebles & Xie, 2023; Ma et al., 2024).
Figure 5: FID-50K over training iterations with different model sizes, where we see clear benefits of scaling up model sizes.
Appendix C Performance vs Training Compute
@@
Figure 6: Comparing the performance vs total training compute comparison of INRFlow and DiT (Peebles & Xie, 2023).
@@
We compare the performance vs total training compute of INRFlow and DiT (Peebles & Xie, 2023) in Gflops. INRFlow-linear denotes the variant of INRFlow where the cross-attention in the spatial aware encoder is replaced with grouping followed by a linear layer. We found this could be an efficient variant of standard INRFlow while still achieving competitive performance. Fig. 6 shows the comparison of the training compute in Gflops vs FID-50K between INRFlow and latent diffusion model DiT (Peebles & Xie, 2023) including the tranining compute of the first stage VAE. We estimate the training cost of VAE based the model card listed in HuggingFace [^1] https://huggingface.co/stabilityai/sd-vae-ft-mse. As shown, the training cost of VAE is not negligible and reasonable models with FID $\approx 6.5$ can be trained for the same cost.
@@
Admittedly, under equivalent training Gflops, INRFlow achieves comparable but not as good performance as DiT in terms of FID score (with a difference smaller than $1.65$ FID points). We attribute this gap to the fact that DiT’s VAE was trained on a dataset much larger than ImageNet, using a domain-specific architecture (e.g. a convolutional U-Net). We believe that the simplicity of implementing and training INRFlow models in practice, and the trivial extension to different data domains (as shown in Sect. F) are strong arguments to counter an FID difference of smaller than $1.65$ points. In addition, applying masking tricks orthogonal to our approach like the ones in (Sehwag et al., 2024) can help mitigate the training compute difference.
In addition, due to the flexibility of cross-attention decoder in INRFlow, one can easily conduct random sub-sampling to reduce the number of decoded coordinate-value pairs during training which ca also saves computation. Fig. 7 shows how number of decoded coordinate-value pairs affects the model performance as well as Gflops in training. An image of resolution 256$\times$256 contains 65536 pixels in total which is the maximal number of coordinate-value pairs during training. As see in Fig. 5(b), a model decoding 4096 coordinate-value pairs saves more than 20% Gflops over one decoding 16384.
Figure 7: FID-50K over training iterations with different number of decoded coordinate-value pairs during training and the corresponding compute cost for a single forward pass.
Appendix D Spatial-aware Latent
Psuedo-coord # latents FID-CLIP($\downarrow$) FID($\downarrow$)
grid 1024 7.32 9.74
random 1024 11.66 17.99
KMeans++ 1024 10.42 15.56
random 2048 8.95 11.86
Table 9: Performance of INRFlow on protein folding.
In image domain, we define pseudo coordinates to lie on a 2D grid, which results in pixels grouping as patches of same size. Tab. 9 lists the results of an ablation study on LSUN-church-256 to compare different pseudo-coordinates. We trained all models for 200K steps with batch size 128 and report results in the table below. INRFlow achieves the best performance when using the default grid pseudo coordinates. When using randomly sampled pseudo coordinates we observe a drop in performance.
We attribute this to the fact that when pseudo-coordinates are randomly sampled, each spatial-aware latent effectively does a different amount of work (since pixels only cross-attend to the nearest pseudo-coordinate). This unbalanced load across latents makes encoding less efficient. There are a few different ways to deal with this without necessarily relying on a grid, one is to cluster similar pseudo-coordinates to provide an equidistant distribution in 2D space (i.e. KMeans++ initialization), another one is to increase the number of spatial-aware latents so that each latent has to do less work. We empirically see that both of this options are effective. Ultimately, having pseudo-coordinates lie on a grid strikes a good balance of efficiency and effectiveness.
Appendix E Architecture Ablation
@@
Table 10: Benchmarking vanilla PerceiverIO and INRFlow with spatially aware latents on LSUN-Church-256 (Yu et al., 2015).
We also provide an architecture ablation in Tab. 10 showcasing different design decisions. We compare two variants of Transformer-based architectures INRFlow: a vanilla PerceiverIO that directly operates on ambient space, but without using spatial aware latents and INRFlow. As it can be seen, the spatially aware latents introduced in INRFlow greatly improve performance across all metrics in the image domain, justifying our design decisions. We note that we did not observe the same large benefits for 3D point clouds, which we hypothesize can be due to their irregular structure.
Appendix F Unconditional 3D Point Cloud generation
Model Layers Hidden size #Latents Heads Decoder layers #Params
INRFlow-B 9 512 1024 4 1 108M
INRFlow-L 12 512 1024 4 1 204M
INRFlow-XL 28 1152 1024 16 1 866M
Table 11: Detailed configurations of INRFlow for point cloud generation.
[…]
Category Model CD EMD CD EMD CD EMD
ShapeGF (Cai et al., 2020) 0.3130 0.6365 45.19 40.25 81.23 80.86
SP-GAN (Li et al., 2021) 0.4035 0.7658 26.42 24.44 94.69 93.95
INRFlow-B (ours) (108M) 0.2861 0.5156 43.38 47.54 75.55 64.95
INRFlow-L (ours) 0.2880 0.5052 44.44 47.16 62.20 62.96
ShapeGF (Cai et al., 2020) 3.7243 2.3944 48.34 44.26 58.01 61.25
SP-GAN (Li et al., 2021) 4.2084 2.6202 40.03 32.93 72.58 83.69
Chair GCA (Zhang et al., 2021) 4.4035 2.5820 45.92 47.89 64.27 64.50
INRFlow-B (ours) (108M) 3.6310 2.1725 46.67 53.31 55.43 51.13
INRFlow-L (ours) 3.5145 2.1860 49.39 49.84 50.52 51.66
ShapeGF (Cai et al., 2020) 1.0200 0.8239 44.03 47.16 61.79 57.24
SP-GAN (Li et al., 2021) 1.1676 1.0211 34.94 31.82 87.36 85.94
Car GCA (Zhang et al., 2021) 1.0744 0.8666 42.05 48.58 70.45 64.20
INRFlow-B (ours) (108M) 0.9923 0.7692 43.46 47.44 60.36 53.27
INRFlow-L (ours) 0.9660 0.7846 44.03 48.86 53.83 54.55
All (55 cat) INRFlow-B (ours) (108M) 3.2586 2.1328 49.00 50.40 54.65 55.70
INRFlow-L (ours) 3.1775 1.9794 49.80 52.39 51.80 53.90
For completeness we also tackle unconditional 3D point cloud generation on ShapeNet (Chang et al., 2015). Note that our model does not require training separate VAEs for point clouds, tuning their corresponding hyper-parameters or designing domain specific networks. We simply adapt our architecture for the change in dimensionality of coordinate-value pairs (e.g. $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ for images to $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ for 3D point clouds.). Note that for 3D point clouds, the coordinates and values are equivalent. In this setting, we compare baselines including LION (Vahdat et al., 2022) which is a recent state-of-the-art approach that models 3D point clouds using a latent diffusion type of approach. Following (Vahdat et al., 2022) we report MMD, COV and 1-NNA as metrics. To have a straightforward comparison with baselines, we train INRFlow-B with to approximately match the number of parameters as LION (Vahdat et al., 2022) (110M for LION vs 108M for INRFlow, see Tab. 11) on the same datasets (using per sample normalization as in Tab. 17 in Vahdat et al. (2022)). On ShapeNet, INRFlow models are trained for 800K iterations with a batch size of 16.
We show results for category specific models and for an unconditional model jointly trained on 55 ShapeNet categories in Tab. 12. INRFlow-B obtains strong generation results on ShapeNet despite being a domain agnostic approach and outperforms LION in most datasets and metrics. Note that INRFlow-B has comparable number of parameters and the same inference settings than LION so this is fair comparison. Finally, we also report results for a larger model INRFlow-L (with ×$\times 2$ the parameter count as LION) to investigate how INRFlow improves as with increasing model size. We observe that with increasing model size, INRFlow typically achieves better performance than the base version. This further demonstrates scalability of our model on ambient space of different data domains. More point cloud samples can be found in Appendix J.
Appendix G Quantitative Results on Protein Folding
Model $C_{\alpha}$-LDDT($\uparrow$) TM-Score($\uparrow$)
Boltz1 (Wohlwend et al., 2024) 0.923 0.812
INRFlow 0.722 0.664
Table 13: Performance of INRFlow on protein folding.
This section includes the quantitative evaluation of the protein folding task. In particular, we randomly selected 512 proteins from the AFDB-SwissProt dataset (Varadi et al., 2022) and use them as a test set. We compare INRFlow with an open-source replication of AlphaFold3 (Abramson et al., 2024) (ie. Boltz-1 (Wohlwend et al., 2024)), which is the SOTA approach for protein folding. Noted that AlphaFold3 is extremely domain-specific, using complex and curated architectural designs for protein folding. For example, it relies on multi-sequence alignment and template search on existing proteins. It also designs a triangle self-attention update for atomic coordinates. Whereas INRFlow makes no assumptions about data domain and effectively models different tasks under an equivalent architecture and training objective. We report $C_{\alpha}$-LDDT and TM-score which are commonly used metrics to evaluate how predicted protein structures align with ground truth (Tab. 13). Results indicate that INRFlow, which uses a domain-agnostic architecture performs decently well on protein folding even when compared to SOTA models that require intricate domain-expertise embedded in the architecture. Note that we have not optimized INRFlow for hyper-parameters in the protein folding experiment.
Appendix H Implementation of Resolution Agnostic Generation
Figure 8: Illustration of resolution agnostic sampling. When generating higher-resolution images in inference (e.g., 512$\times$512), 256$\times$256 coordinate-value pairs (consistent to setting in training) are selected through grid subsampling and are fed to the encoder. The decoder takes in the full 512$\times$512 coordinate-value pairs to predict velocities of all the pixel values. The model repeats the process in inference to generate a 512$\times$512 image.
One can interpret the spatial-aware latents computed from INRFlow’s encoder as the “latent codes that are transformed into network parameters”. For INRFlow these latents codes are used to compute $v,k$ for the cross-attention block in the decoder, which takes in queries $q$ (i.e. coordinate-value pairs) at any resolution. The only thing that we need to do in order to obtain consistent outputs across different query resolutions is to keep the resolution of the encoder fixed, while the decoder can be queried at a different resolution. This is trivially achieved by employing a simple sub-sampling operation (i.e. grid sub-sampling) and keeping the sub-sampling operation fixed during inferenc (Fig. 8). This simple technique allows us to change the resolution at inference time without any other additional tricks regarding noise alignment at different resolutions and produces crisp and consistent examples at higher resolutions that the one used in training (more results see Fig. 4 and 14).
Appendix I Additional ImageNet Samples
@@
We show uncurated samples of different classes from INRFlow-XL trained on ImageNet-256 in Fig. 9 and Fig. 10. Guidance scales in CFG are set as $4.0$ for loggerhead turtle, macaw, otter, coral reef and $2.0$ otherwise.
@@
Figure 9: Uncurated samples of class labels: loggerhead turtle (33), macaw (88), golden retriever (207), otter (360) and red panda (387), and panda (388) from INRFlow trained on ImageNet-256.
@@
Figure 10: Uncurated samples of class labels: palace (698), space shuttle (812), ice cream (928), pizza (963), coral reef (973), and valley (979) from INRFlow trained on ImageNet-256.
Appendix J Additional ShapeNet Samples
We show uncurated samples from INRFlow-L trained jointly on 55 ShapeNet categories in Fig. 11.
@@
Appendix K Additional Objaverse Samples
We show additional Objaverse samples from INRFlow in Fig. 12
Figure 12: Image-conditioned point clouds with 16k, 32k, 64k, and 128k points generated from an INRFlow trained on Objaverse (training with 16k points, CFG scale 5.0).
Appendix L Additional SwissProt Samples
We show additional SwissProt samples from INRFlow in Fig. 13
Figure 13: Additional examples of protein structures predicted by INRFlow on SwissProt (Boeckmann et al., 2003)
Appendix M Additional Resolution Agnostic Image Samples
We show additional samples generated at different resolutions from INRFlow trained on ImageNet-256 in Fig. 14.
Figure 14: Images generated at 256, 512, 1024, and 2048 resolutions from an INRFlow trained on ImageNet-256.